Jefferson Steel requires $15 million to fund its current year’s capital projects. Jefferson will finance part of its needs with equity. The firm’s common stock is selling in the market at $180 per share. Dividends of $1 per share were recently paid (Do). Dividend growth of 6 percent per year is expected for the foreseeable future. The market is currently demanding a 6 percent premium on the average risk stock and Tbonds are currently yielding 3%. Preferred stock is selling in the market at $97 per share. The preferred has a dividend rate of 4.5% percent on $200 face value per share. The remainder of needs will be financed with debt. Jefferson's current bonds are ten year, $1000 par bonds with a coupon rate of 8% and are selling at a price of $1020 per bond. Interest is paid semi-annually on the bonds. The firm faces a 30 percent marginal tax rate and the project asset beta is estimated at 1.45. Jefferson believes the capital asset pricing model is the best way to estimate the cost of equity. Jefferson’s balance sheet, in part, is presented below. The number of units outstanding are shown in parenthesis.

($ x millions) Accounts Payable Accruals LT Debt ($1000 par) Total Debt $17,000,000 Pfd Stock 14,000,000 Common stock @ par (What is the cost of debt to the firm? Rd =

Estimate the cost of preferred stock to the firm (Rp) Rp =

Estimate the cost of equity to the firm (Re) Re =
Answer #1

Rd = YTM of bonds * (1 - tax rate)

YTM is calculated using RATE function in Excel with these inputs :

nper = 10*2 (10 years to maturity with 2 semiannual coupon payments each year)

pmt = 1000 * 8% / 2 (semiannual coupon payment = face value * annual coupon rate / 2. This is a positive figure as it is an inflow to the bondholder)

pv = -1020 (current bond price. This is a negative figure as it is an outflow to the buyer of the bond)

fv = 1000 (face value of the bond receivable on maturity. This is a positive figure as it is an inflow to the bondholder)

the RATE is calculated to be 3.85%. This is the semiannual YTM. To calculate the annual YTM, we multiply by 2. Annual YTM is 7.71%

A1 fx =RATE(10*2,1000*8%/2,-1020, 1000) D E F B C A 11 3.85%

cost of debt = YTM * (1 - tax rate)

cost of debt = 7.71% * (1 - 30%) ==> 5.40%

Rd = 5.40%

Rp = annual dividend amount / price per share

annual dividend amount = face value * dividend rate = $200 * 4.5% = $9.

Rp = $9 / $97

Rp = 9.28%

Re = risk free rate + (beta * market risk premium)

Re = 3% + (1.45 * 6%)

Re = 11.7%